# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ V Datasets and Evaluation
V-A Training Datasets
The performance of LLMs largely depends on the training data’s quality, size, and diversity. Preparing training datasets of high quality at a large scale is laborious. Researchers have suggested various pre-training and fine-tuning datasets to enhance LLMs capabilities. We summarize these efforts in Table VIII. While numerous training datasets are available in the literature, we cover the most widely used ones in our summary.
V-B Evaluation Datasets and Tasks
@@
Numerous datasets are proposed for each task, evaluating LLMs against different characteristics. To provide an overview of evaluation datasets, we briefly discuss a few famous datasets within each category and offer a comprehensive list of datasets in Table IX. Moreover, we show a detailed overview of the training datasets and evaluation tasks and benchmarks used by various pre-trained LLMs in Table X and fine-tuned LLMs in Table XI. We also compare the top-performing LLMs in various NLP tasks in Table XII.
TABLE VIII: Details of various well-known pre-training and fine-tuning datasets. Here, alignment means aligning with human preferences.
Dataset Type Size/Samples Tasks Source Creation
Comments
C4 [10] Pretrain 806GB - Common Crawl Automated
A clean, multilingual dataset with billions of tokens
mC4 [11] Pretrain 38.49TB - Common Crawl Automated
A multilingual extension of the C4 dataset, mC4 identifies over 100 languages using cld3 from 71 monthly web scrapes of Common Crawl.
@@
Automated
A massive dataset comprised of 22 constituent sub-datasets
ROOTs [291] Pretrain 1.61TB - 498 Hugging Face datasets Automated
46 natural and 13 programming languages
@@
Automated
99% of the data is in English
Wikipedia [292] Pretrain - - Wikipedia Automated
Dump of wikipedia
@@
Automated
Open-source replica of LLaMA dataset
PushShift.io Reddit Pretrain 21.1GB - Reddit Automated
Submissions and comments on Reddit from 2005 to 2019
BigPython [129] Pretrain 5.5TB Coding GitHub Automated
Pool of Prompt (P3) [17] Instructions 12M 62 PromptSource Manual
A Subset of PromptSource, created from 177 datasets including summarization, QA, classification, etc.
xP3 [143] Instructions 81M 71 P3+Multilingual datasets Manual
Extending P3 to total 46 languages
Super-NaturalInstructions (SNI) [18] Instructions 12.4M 1616 Multiple datasets Manual
Extending P3 with additional multi-lingual datasets, total 46 languages
Flan [16] Instructions 15M 1836 Muffin+T0-SF+NIV2 Manual
Total 60 languages
OPT-IML [92] Instructions 18.1M 1667 - Manual
Self-Instruct [19] Instructions 82k 175 - Automated
Generated 52k instructions with 82k samples from 175 seed tasks using GPT-3
Alpaca [147] Instructions 52k - - Automated
Employed self-instruct method to generate data from text-davinci-003
Vicuna [148] Instructions 125k - ShareGPT Automated
Conversations shared by users on ShareGPT using public APIs
LLaMA-GPT-4 [149] Instructions 52k - Alpaca Automated
Recreated Alpaca dataset with GPT-4 in English and Chinese
Unnatural Instructions [294] Instructions 68k - 15-Seeds (SNI) Automated
LIMA [174] Instructions 1k - Multiple datasets Manual
Carefully created samples to test performance with fine-tuning on less data
Anthropic-HH-RLHF [295] Alignment 142k - - Manual
Anthropic-HH-RLHF-2 [167] Alignment 39k - - Manual
TABLE IX: Categorized evaluation datasets used in evaluating LLMs.
@@
Multi-Task
MMLU [296], SuperGLUE [2], BIG-bench [297], GLUE [298], BBH [297], CUGE [299], ZeroCLUE [300], FewCLUE [301], Blended Skill Talk [302], HELM [303], KLUE-STS [304]
Language Understanding
CoQA [305], WiC [306], Wikitext103 [307], PG19 [308], LCQMC [309], QQP [310], WinoGender [311], CB [312], FinRE [313], SanWen [314], AFQMC [300], BQ Corpus [315], CNSS [316], CKBQA 13 [317], CLUENER [300], Weibo [318], AQuA [319], OntoNotes [320], HeadQA [321], Twitter Dataset [322]
@@
RACE [336], RACE-Middle [336], RACE-High [336], QuAC [337], StrategyQA [338], Quiz Bowl [339],
cMedQA [340], cMedQA2 [341], MATINF-QA [342]
Commonsense Reasoning
WinoGrande [343], HellaSwag [344], COPA [345], WSC [346], CSQA [347], SIQA [348], C3 [349], CLUEWSC2020 [300], CLUEWSC [300], CLUEWSC-FC [301], ReCoRD [350]
Reading Comprehension
SQuAD [351], BoolQ [352], SQUADv2 [353], DROP [354], RTE [355], WebQA [356], CMRC2017 [357], CMRC2018 [358], CMRC2019 [359], COTE-BD [360], COTE-DP [360], COTE-MFW [360], MultiRC [361], Natural Questions [362], CNSE [316], DRCD [363], DuReader [364], Dureaderrobust [365], DuReader-QG [364], SciQ [366], Sogou-log [367], Dureaderrobust-QG [365], QA4MRE [368], KorQuAD 1.0 [369], CAIL2018-Task1 & Task2 [370]
Mathematical Reasoning
MATH [371], Math23k [372], GSM8K [373], MathQA [374], MGSM [375], MultiArith [376], ASDiv [377], MAWPS [378], SVAMP [379]
Problem Solving
HumanEval [130], DS-1000 [380], MBPP [381], APPS [371], CodeContests [131]
@@
ANLI [382], MNLI-m [383], MNLI-mm [383],QNLI [351], WNLI [346], OCNLI [300], CMNLI [300], ANLI R1 [382], ANLI R2 [382], ANLI R3 [382], HANS [384], OCNLI-FC [301], LogiQA [385], StrategyQA [338]
Cross-Lingual Understanding
MLQA [386], XNLI [387], PAWS-X [388], XSum [389], XCOPA [390], XWinograd [391], TyDiQA-GoldP [392], MLSum [393]
Truthfulness and Fact Checking
TruthfulQA [394], MultiFC [395], Fact Checking on Fever [396]
Biases and Ethics in AI
ETHOS [397], StereoSet [398], BBQ [399], Winobias [400], CrowS-Pairs [401]
Toxicity
RealToxicityPrompts [402], CivilComments toxicity classification [403]
Language Translation
WMT [404], WMT20 [405], WMT20-enzh [405], EPRSTMT [301], CCPM [406]
Scientific Knowledge
AminoProbe [137], BioLAMA [137], Chemical Reactions [137], Galaxy Clusters [137], Mineral Groups [137]
Dialogue
Wizard of Wikipedia [407], Empathetic Dialogues [408], DPC-generated [119] dialogues, ConvAI2 [409], KdConv [410]
Topic Classification
TNEWS-FC [301], YNAT [304], KLUE-TC [304], CSL [300], CSL-FC [301], IFLYTEK [411]
TABLE X: An illustration of training datasets and evaluation tasks employed by pre-trained LLMs. Here, “QA” is question-answering, “Clf” is classification, “NLI” is natural language inference, “MT” is machine translation, “RC” is reading comprehension, “CR” is commonsense reasoning, “MR” is mathematical reasoning, “Mem.” is memorization.
Benchmark
Models
Training Dataset
@@
T5
C4 [10]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-3
Common Crawl, WebText, Books Corpora, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
mT5
mC4 [11]
$\checkmark$ $\checkmark$ $\checkmark$
PanGu-$\alpha$
1.1TB Chinese Text Corpus
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CPM-2
WuDaoCorpus [104]
$\checkmark$ $\checkmark$
Codex
54 million public repositories from Github
$\checkmark$
ERNIE-3.0
@@
Chinese text corpora, Baidu Search, Web text, QA-long, QA-short, Poetry and Couplet Domain-specific data from medical, law, and financial area Baidu knowledge graph with more than 50 million facts
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Jurassic-1
Wikipedia, OWT, Books, C4, Pile [290], arXiv, GitHub
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
HyperCLOVA
@@
Korean blogs, Community sites, News, KiN Korean Wikipedia, Wikipedia (English and Japanese), Modu-Corpus: Messenger, News, Spoken and written language corpus, Web corpus
$\checkmark$
Yuan 1.0
@@
Common Crawl, SogouT, Sogou News, Baidu Baike, Wikipedia, Books
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Gopher
@@
subsets of MassiveWeb Books, C4, News, GitHub and Wikipedia samples from MassiveText
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ERNIE-3.0 TITAN
Same as ERNIE 3.0 and ERNIE 3.0 adversarial dataset, ERNIE 3.0 controllable dataset
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-NeoX-20B
Pile [290]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPT
RoBERTa [288], Pile [290], PushShift.io Reddit [412]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BLOOM
ROOTs [13]
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Galactica
@@
arXiv, PMC, Semantic Scholar, Wikipedia, StackExchange, LibreText, Open Textbooks, RefSeq Genome, OEIS, LIPID MAPS, NASAExoplanet, Common Crawl, ScientificCC, AcademicCC, GitHub repositories Khan Problems, GSM8K, OneSmallStep
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GLaM
@@
Filtered Webpages, Social media conversations Wikipedia, Forums, Books, News
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LaMDA
Infiniset : Public documents, Dialogs, Utterances
$\checkmark$
MT-NLG
@@
Two snapshots of Common Crawl and Books3, OpenWebText2, Stack Exchange, PubMed Abstracts, Wikipedia, PG-19 [242], BookCorpus2, NIH ExPorter, Pile, CC-Stories, RealNews
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AlphaCode
@@
Selected GitHub repositories, CodeContests: Codeforces, Description2Code, CodeNet
$\checkmark$
Chinchilla
@@
MassiveWeb, MassiveText Books, C4, News, GitHub, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PaLM
@@
webpages, books, Wikipedia, news, articles, source code, social media conversations
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AlexaTM
Wikipedia, mC4
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
U-PaLM
Same as PaLM
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
UL2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GLM-130B
$\checkmark$ $\checkmark$ $\checkmark$
CodeGen
Pile, BigQuery, BigPython
$\checkmark$
LLaMA
@@
CommonCrawl, C4, Github, Wikipedia, Books, arXiv, StackExchange
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PanGu-$\Sigma$
@@
WuDaoCorpora, CLUE, Pile, C4, Python code
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BloombergGPT
inPile, Pile, C4, Wikipedia
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CodeT5+
CodeSearchNet, Github Code
$\checkmark$ $\checkmark$
StarCoder
The Stack v1.2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LLaMA-2
$\checkmark$
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PaLM-2
Web documents, Code, Books, Maths, Conversation
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
TABLE XI: An illustration of training datasets and evaluation benchmarks used in fine-tuned LLMs. “SNI” is a short of Super-NaturalInsturctions.
Models
Training Dataset
@@
T0
Pool of Prompts
$\checkmark$
WebGPT
@@
ELI5 [413], ELI5 fact-check [155], TriviaQA [330], ARC-Challenge [331], ARC-Easy [331], Hand-written data, Demonstrations of humans, Comparisons between model-generated answers
$\checkmark$
Tk-INSTRUCT
SNI [18]
$\checkmark$
mT0
xP3 [143]
OPT-IML
PromptSource [17], FLAN [16], SNI [414], UnifiedSKG [415], CrossFit [416], ExMix [417], T5 [10], Reasoning
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Flan
Muffin, T0-SF, NIv2, CoT
$\checkmark$ $\checkmark$ $\checkmark$
WizardCoder
Code Alpaca
$\checkmark$ $\checkmark$
V-B1 Multi-task
MMLU [296]

### 09 Apr 2024 v9

@@ 5 Datasets and Evaluation
5.1 Training Datasets
The performance of LLMs largely depends on the training data’s quality, size, and diversity. Preparing training datasets of high quality at a large scale is laborious. Researchers have suggested various pre-training and fine-tuning datasets to enhance LLMs capabilities. We summarize these efforts in Table 8. While numerous training datasets are available in the literature, we cover the most widely used ones in our summary.
5.2 Evaluation Datasets and Tasks
@@
Numerous datasets are proposed for each task, evaluating LLMs against different characteristics. To provide an overview of evaluation datasets, we briefly discuss a few famous datasets within each category and offer a comprehensive list of datasets in Table 9. Moreover, we show a detailed overview of the training datasets and evaluation tasks and benchmarks used by various pre-trained LLMs in Table 10 and fine-tuned LLMs in Table 11. We also compare the top-performing LLMs in various NLP tasks in Table 12.
Table 8: Details of various well-known pre-training and fine-tuning datasets. Here, alignment means aligning with human preferences.
[…]
C4 [10] Pretrain 806GB - Common Crawl Automated A clean, multilingual dataset with billions of tokens
mC4 [11] Pretrain 38.49TB - Common Crawl Automated A multilingual extension of the C4 dataset, mC4 identifies over 100 languages using cld3 from 71 monthly web scrapes of Common Crawl.
@@
Automated A massive dataset comprised of 22 constituent sub-datasets
ROOTs [292] Pretrain 1.61TB - 498 Hugging Face datasets Automated 46 natural and 13 programming languages
@@
Automated 99% of the data is in English
Wikipedia [293] Pretrain - - Wikipedia Automated Dump of wikipedia
@@
Automated Open-source replica of LLaMA dataset
PushShift.io Reddit Pretrain 21.1GB - Reddit Automated Submissions and comments on Reddit from 2005 to 2019
BigPython [130] Pretrain 5.5TB Coding GitHub Automated -
Pool of Prompt (P3) [17] Instructions 12M 62 PromptSource Manual A Subset of PromptSource, created from 177 datasets including summarization, QA, classification, etc.
xP3 [144] Instructions 81M 71 P3+Multilingual datasets Manual Extending P3 to total 46 languages
Super-NaturalInstructions (SNI) [18] Instructions 12.4M 1616 Multiple datasets Manual Extending P3 with additional multi-lingual datasets, total 46 languages
Flan [16] Instructions 15M 1836 Muffin+T0-SF+NIV2 Manual Total 60 languages
OPT-IML [97] Instructions 18.1M 1667 - Manual -
Self-Instruct [19] Instructions 82k 175 - Automated Generated 52k instructions with 82k samples from 175 seed tasks using GPT-3
Alpaca [148] Instructions 52k - - Automated Employed self-instruct method to generate data from text-davinci-003
Vicuna [149] Instructions 125k - ShareGPT Automated Conversations shared by users on ShareGPT using public APIs
LLaMA-GPT-4 [150] Instructions 52k - Alpaca Automated Recreated Alpaca dataset with GPT-4 in English and Chinese
Unnatural Instructions [295] Instructions 68k - 15-Seeds (SNI) Automated -
LIMA [175] Instructions 1k - Multiple datasets Manual Carefully created samples to test performance with fine-tuning on less data
Anthropic-$H$-RLHF [296] Alignment 142k - - Manual
Anthropic-$H$-RLHF-2 [168] Alignment 39k - - Manual
Table 9: Categorized evaluation datasets used in evaluating LLMs.
@@
Multi-Task MMLU [297], SuperGLUE [2], BIG-bench [298], GLUE [299], BBH [298], CUGE [300], ZeroCLUE [301], FewCLUE [302], Blended Skill Talk [303], HELM [304], KLUE-STS [305]
Language Understanding CoQA [306], WiC [307], Wikitext103 [308], PG19 [309], LCQMC [310], QQP [311], WinoGender [312], CB [313], FinRE [314], SanWen [315], AFQMC [301], BQ Corpus [316], CNSS [317], CKBQA 13 [318], CLUENER [301], Weibo [319], AQuA [320], OntoNotes [321], HeadQA [322], Twitter Dataset [323]
@@
RACE [337], RACE-Middle [337], RACE-High [337], QuAC [338], StrategyQA [339], Quiz Bowl [340], cMedQA [341],cMedQA2 [342], MATINF-QA [343]
Commonsense Reasoning WinoGrande [344], HellaSwag [345], COPA [346], WSC [347], CSQA [348], SIQA [349], C3 [350], CLUEWSC2020 [301], CLUEWSC [301], CLUEWSC-FC [302], ReCoRD [351]
Reading Comprehension SQuAD [352], BoolQ [353], SQUADv2 [354], DROP [355], RTE [356], WebQA [357], CMRC2017 [358], CMRC2018 [359], CMRC2019 [360], COTE-BD [361], COTE-DP [361], COTE-MFW [361], MultiRC [362], Natural Questions [363], CNSE [317], DRCD [364], DuReader [365], Dureaderrobust [366], DuReader-QG [365], SciQ [367], Sogou-log [368], Dureaderrobust-QG [366], QA4MRE [369], KorQuAD 1.0 [370], CAIL2018-Task1 & Task2 [371]
Mathematical Reasoning MATH [372], Math23k [373], GSM8K [374], MathQA [375], MGSM [376], MultiArith [377], ASDiv [378], MAWPS [379], SVAMP [380]
Problem Solving HumanEval [131], DS-1000 [381], MBPP [382], APPS [372], CodeContests [132]
@@
ANLI [383], MNLI-m [384], MNLI-$m$ [384],QNLI [352], WNLI [347], OCNLI [301], CMNLI [301], ANLI R1 [383], ANLI R2 [383], ANLI R3 [383], HANS [385], OCNLI-FC [302], LogiQA [386], StrategyQA [339]
Cross-Lingual Understanding MLQA [387], XNLI [388], PAWS-X [389], XSum [390], XCOPA [391], XWinograd [392], TyDiQA-GoldP [393], MLSum [394]
Truthfulness and Fact Checking TruthfulQA [395], MultiFC [396], Fact Checking on Fever [397]
Biases and Ethics in AI ETHOS [398], StereoSet [399], BBQ [400], Winobias [401], CrowS-Pairs [402]
Toxicity RealToxicityPrompts [403], CivilComments toxicity classification [404]
Language Translation WMT [405], WMT20 [406], WMT20-enzh [406], EPRSTMT [302], CCPM [407]
Scientific Knowledge AminoProbe [138], BioLAMA [138], Chemical Reactions [138], Galaxy Clusters [138], Mineral Groups [138]
Dialogue Wizard of Wikipedia [408], Empathetic Dialogues [409], DPC-generated [96] dialogues, ConvAI2 [410], KdConv [411]
Topic Classification TNEWS-FC [302], YNAT [305], KLUE-TC [305], CSL [301], CSL-FC [302], IFLYTEK [412]
Table 10: An illustration of training datasets and evaluation tasks employed by pre-trained LLMs. Here, “QA” is question-answering, “Clf” is classification, “NLI” is natural language inference, “MT” is machine translation, “RC” is reading comprehension, “CR” is commonsense reasoning, “MR” is mathematical reasoning, “Mem.” is memorization.
[…]
@@
T5 C4 [10] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-3 Common Crawl, WebText, Books Corpora, Wikipedia $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
mT5 mC4 [11] $\checkmark$ $\checkmark$ $\checkmark$
PanGu-$\alpha$ 1.1TB Chinese Text Corpus $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CPM-2 WuDaoCorpus [109] $\checkmark$ $\checkmark$
Codex 54 million public repositories from Github $\checkmark$
@@
ERNIE-3.0 Chinese text corpora, Baidu Search, Web text, QA-long, QA-short, Poetry and Couplet Domain-specific data from medical, law, and financial area Baidu knowledge graph with more than 50 million facts $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Jurassic-1 Wikipedia, OWT, Books, C4, Pile [291], arXiv, GitHub $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
HyperCLOVA Korean blogs, Community sites, News, KiN Korean Wikipedia, Wikipedia (English and Japanese), Modu-Corpus: Messenger, News, Spoken and written language corpus, Web corpus $\checkmark$
@@
Yuan 1.0 Common Crawl, SogouT, Sogou News, Baidu Baike, Wikipedia, Books $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Gopher subsets of MassiveWeb Books, C4, News, GitHub and Wikipedia samples from MassiveText $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ERNIE-3.0 TITAN Same as ERNIE 3.0 and ERNIE 3.0 adversarial dataset, ERNIE 3.0 controllable dataset $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GPT-NeoX-20B Pile [291] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPT RoBERTa [289], Pile [291], PushShift.io Reddit [413] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BLOOM ROOTs [13] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Galactica arXiv, PMC, Semantic Scholar, Wikipedia, StackExchange, LibreText, Open Textbooks, RefSeq Genome, OEIS, LIPID MAPS, NASAExoplanet, Common Crawl, ScientificCC, AcademicCC, GitHub repositories Khan Problems, GSM8K, OneSmallStep $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
GLaM Filtered Webpages, Social media conversations Wikipedia, Forums, Books, News $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LaMDA Infiniset : Public documents, Dialogs, Utterances $\checkmark$
@@
MT-NLG Two snapshots of Common Crawl and Books3, OpenWebText2, Stack Exchange, PubMed Abstracts, Wikipedia, PG-19 [242], BookCorpus2, NIH ExPorter, Pile, CC-Stories, RealNews $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
AlphaCode Selected GitHub repositories, CodeContests: Codeforces, Description2Code, CodeNet $\checkmark$
@@
Chinchilla MassiveWeb, MassiveText Books, C4, News, GitHub, Wikipedia $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
PaLM webpages, books, Wikipedia, news, articles, source code, social media conversations $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AlexaTM Wikipedia, mC4 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
U-PaLM Same as PaLM $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
UL2 - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
GLM-130B - $\checkmark$ $\checkmark$ $\checkmark$
CodeGen Pile, BigQuery, BigPython $\checkmark$
@@
LLaMA CommonCrawl, C4, Github, Wikipedia, Books, arXiv, StackExchange $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
PanGu-$\Sigma$ WuDaoCorpora, CLUE, Pile, C4, Python code $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
BloombergGPT inPile, Pile, C4, Wikipedia $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CodeT5+ CodeSearchNet, Github Code $\checkmark$ $\checkmark$
StarCoder The Stack v1.2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
LLaMA-2 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PaLM-2 Web documents, Code, Books, Maths, Conversation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Table 11: An illustration of training datasets and evaluation benchmarks used in fine-tuned LLMs. “SNI” is a short of Super-NaturalInsturctions.
[…]
@@
T0 Pool of Prompts $\checkmark$
@@
WebGPT ELI5 [414], ELI5 fact-check [156], TriviaQA [331], ARC-Challenge [332], ARC-Easy [332], Hand-written data, Demonstrations of humans, Comparisons between model-generated answers $\checkmark$
Tk-INSTRUCT SNI [18] $\checkmark$
mT0 xP3 [144]
OPT-IML PromptSource [17], FLAN [16], SNI [415], UnifiedSKG [416], CrossFit [417], ExMix [418], T5 [10], Reasoning $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Flan Muffin, T0-SF, NIv2, CoT $\checkmark$ $\checkmark$ $\checkmark$
WizardCoder Code Alpaca $\checkmark$ $\checkmark$
5.2.1 Multi-task